OpenStudy (anonymous):

C(66;25) / C(70;25) How would I solve this? ^

OpenStudy (anonymous):

I just need it interpreted mainly

OpenStudy (anonymous):

@Luis_Rivera @Mertsj

OpenStudy (mertsj):

Is that 66 choose 25 divided by 70 choose 25 or what?

OpenStudy (anonymous):

it has to do with probability

OpenStudy (anonymous):

the semicolon can be a comma too

OpenStudy (mertsj):

\[\frac{66!}{25!(66-25)!}\]

OpenStudy (mertsj):

Evaluate that and divide it by this: \[\frac{70!}{25!(70-25!)}\]

OpenStudy (anonymous):

im getting an error for the 2nd one

OpenStudy (anonymous):

got 2.933706409 for the first

OpenStudy (anonymous):

my calc says: work overflow

OpenStudy (mertsj):

\[\frac{\frac{66!}{25!41!}}{\frac{70!}{25!45!}}=\frac{66!}{25!41!}\times \frac{25!45!}{70!}=\]

OpenStudy (mertsj):

\[\frac{66!45\times44\times43\times42\times41!}{41!70\times69\times68\times67\times66!}=\]

OpenStudy (mertsj):

\[\frac{45\times44\times 43\times 42}{70\times69\times68\times67}\]

OpenStudy (mertsj):

Perhaps you can finish it.

OpenStudy (anonymous):

What is the probability that at least 2 people in a random group of 6 people have a birthday on the same day of the week?

OpenStudy (anonymous):

can you help me with this question too?

OpenStudy (anonymous):

the answer is p(7,6)/(7^6)=.0428 1-.0428=9572.... but i dont know how to put the numerator into a calc.

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

p(7,6)=7!/(7-6)!

OpenStudy (mertsj):